OpenStudy (anonymous):

The roof lines of a building can be described by the system of equations below, where the floor is represented by the x-axis and the y-axis and the height of the building is along the z-axis. Measurements are in feet. Find the point of intersection of the roof lines, using an ordered triple in the form (x, y, z). x+y+z=53 3x-2y+z=69 -x+2y-z=-59

Miracrown (miracrown):

ok, so we need to solve for (x,y,z) we could use substitution, or a matrix do you have a preference ?

OpenStudy (kainui):

This one's pretty nice, you can add the first equation to the last equation and both the x and z terms cancel out.

OpenStudy (anonymous):

I figured it out !!!

OpenStudy (jack1):

\(\large x+y+z=53 ~~~~\rightarrow ENQ~1\) so: \(\large x = 53-y-z~~~~ \rightarrow ENQ~1a\) \(\large -x+2y-z=-59 ~~~~\rightarrow ENQ~3\) sub EQN 1a into EQN 3 \(\large -x+2y-z=-59 \) \(\large -(53-y-z)+2y-z=-59 \) \(\large -53+y+z+2y-z=-59 \) \(\large 3y=-59+53 \) \(\Large y=\frac {-6}3 = -2 \) now you know y, just continue from here with substitution until you know x and z the intersection point will be at the 3 dimensional point (x,y,z)